OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

Plug in 3 for n, then 4, then 5.\[f(3)=f(1)+f(2)+f(3-1)=f(1)+f(2)+f(2)=2+4+4=10\]\[f(4))=f(1)+f(2)+f(4-1)=f(1)+f(2)+f(3)=2+4+10=16\]\[f(5)=f(1)+f(2)+f(5-1)=f(1)+f(2)+f(4)=2+4+16=22\]So f(5)=22

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

No problem, you're welcome